OpenStudy (anonymous):

Evaluate the expression below. Write your answer to three decimal places. log₆95

OpenStudy (mathstudent55):

You need change of base formula since your calculator can do log base 10 and maybe log base e, but not log of any other base.

OpenStudy (mathstudent55):

\[\log _{a} x = \frac{ \log _{b} x }{ \log _{b} a}\]

OpenStudy (mathstudent55):

This means to take the log (unusual base) of a number, take log (base 10 or base e) of the number and divide by log (base 10 or base e) of unusual base.

OpenStudy (mathstudent55):

\[\log _{6} 95 = \frac{ \log _{10} 95 }{ \log _{10} 6}\]

OpenStudy (anonymous):

thank you so much!! so is that the final answer?

OpenStudy (anonymous):

i'll go over it more so i understand better

OpenStudy (mathstudent55):

2.542

OpenStudy (anonymous):

you are amazing, thanks so so much